ACIS SIM Integration Test Page

Latest update: Tue Feb 3 10:52:41 1998 EST

ACIS Home Page Status Plan Events E-Notes S-Notes Anomalies Phone Numbers
Live Engineering Data Procedures
Long Form Functional Short Form functional Test Matrix

Current Status:

  • Started pump down of BRUTUS 7pm local, Wednesday, 1/7
  • Started Ambient Vacuum SIM Functional 11am local, Thursday, 1/8
  • Started first cold->hot transition 10pm, Saturday, 1/10
  • Started second cold->hot transition 11pm, Monday, 1/12
  • Started Thermal Balance soak, 0010, Thursday, 1/15
  • Started Eclipse simulation 0045, Saturday, 1/17
  • Broke vacuum 1500, Sunday, 1/18
  • Pre-ship Functional 1500, Saturday, 1/25
  • Ship to TRW, Friday, 1/30

    Short Term Plan:


    Test Event History

    Day Planned Event Actual Events Results
    7/2 Arrive at BASD Did that
    7/10 Long Form DEA/DPA/SS only OK
    7/14 Flight Software Burn Version 1.5 loaded BEP A had no errors
    BEP B did not program
    7/15 Test FS Found GSE problem which precluded burning BEP B yesterday Flight 1.5 testing OK so far.
    7/16 Test FS; burn BEP B Did that BEP B programmed without problem
    Scheduled testing completed
    7/23 CTUE/RCTU BER Test Ran with XRCF Cables 10**7 bits flowed w/o errors.
    7/24 CTUE/RCTU BER Test Ran with Flight Cables 2*10**7 bits flowed w/o errors using new tlm format.
    7/30 RCTU<->ACIS Signal Verification Did active/passive telemetry; pulse commands All per plan.
    7/31 RCTU<->ACIS Signal Verification Did Serial Telemetry, A and B All per plan
    8/1 RCTU<->ACIS Signal Verification Did Serial Commands A and B Leading edges ringing like a doorbell
    Timing is OK.
    8/2 RCTU<->ACIS Signal Verification Looked at Serial Cmd Clock at DPA. Rings here, too.
    8/5 RCTU<->ACIS Signal Verification More ways to look at Serial Cmd Clock. Ambiguous results.
    8/12 RCTU<->ACIS Signal Verification Looked at many things. Ambiguous results.
    9/16 PSMC Mounted on ISIM Mechanics OK. Only power cables didn't fit.
    9/20 Detector Assembly onto SIM It went after interferences fixed. Vent Valve shims badly done.
    9/21 Support Structure onto SIM
    Went well.
    9/22 ACIS Short Form Even ran DEA B. Noise at the 15ADU level.
    9/24 ISIM Alignment Mostly done on 9/25 MWB to analyze results
    9/26 ISIM Crosstalk Tests Completed 1900, 9/27 No noise noted.
    10/10 ISIM Installed in Brutus
    HST miswiring being worked
    10/12 Run pre-TB Long Form
    Looks OK; HST OK
    10/19 Pump Doun Door Closed at 12:45 Pumping started at 13:00
    10/20 Ambient Short Form Done at 16:30 Nominal results
    10/20 Open Big Valve and Door Valve at 1715; door at 2300 About 1.5torr in box
    10/21 Start 6 CCD ops (ACIS in focus) 36 hr stabilization officially started at 0400 local Focal plane stable at -110C
    10/22 Start S2 at HRC-S Focus Started at 1745 local in 500bps mode (for 6 hr) Guesses on windows/flux very good.
    10/23 Start S4 at HRC-S Focus Phase 3a started at 1030 local
    10/24 Start Cold Operating/2 CCD Phase 5 started at 0030 local NB we skipped 3b, 4 to get here.
    10/25 Eclipse Mode & ACIS Standby Eclipse started 1000; Standby started 1230
    10/25 Power off for Cold Survival Cold Survival started 1400 Focal Plane hit -129C with radiator at -135C
    Detector Assy hit -75C with radiator at -92C
    10/27 S2 at HRC-I position Started 0800
    10/27 Crosstalk Test Finished with HRC on 10/28
    10/28 Bake out and Close up DA Finished 0100 on 10/29
    10/30 Break Vacuum 2hr repress finished at 0600 Took off PSMC for capacitor changeout.
    11/14 Repress DA and check OBF Both filters look good Also take off shades for rework.
    12/31 Put ISIM back in BRUTUS
    HST verified on 1/5
    1/7/98 Start pumpdown for TV Start time was 1900 local
    1/9 Closed all DA valves Valves closed at 0900 Chamber cool down started 0200 local
    1/11 Run SF during cold->hot Completed at 0500
    1/12 Run science during cold->hot LED, CC, Graded Histogram OBF looks OK
    1/13 Open Large Valve beginning of hot->cold Getting ready for TB soak
    1/14 Cold survival exposure 12hr soak; fairly stable after 6hr Focal Plane hit -128C with radiator -135/-133
    Detector Assy hit -72.2 with radiator -89/-92
    1/15 Thermal Balance Cold retest Run door source science; retest Graded Histogram Focal Plane stable at -122C with radiator -128C;
    Detector Assy controlled to -60C.
    1/17 Eclipse Simulation Followed by bakeout; valve closing
    1/18 Break Vacuum 1500 MST
    1/24 Pre-Ship Functional Normal


    Test Notes from Engineering

    • 7/14: One LRCTU isn't fully functional; we're shipping it home.
    • 7/23: EGSE -> CTUE connection will not support commands in unlimited haste.
      (The BER test needs the 1 second sleeps.)
    • 7/24: Scott Texter and Mike Doucette worked for many hours in the clean room trying to connect the CTUE to the Flilght RCTU. Turns out there is a wiring error in the Flight Cable set. Though the O-ring was changed on the DA GSE pressure transducer, the leak continues at the same old rate.
      The BER test ran fine with the flight cables and the new format, though we know Format 1 won't work for science, and Format 2 isn't reporting bi-levels yet.
    • 7/31: Scott Texter and Mike Zimmerman (Gulton) and Mike Doucette have worked long hours in the clean room doing RCTU<->ACIS electrical signal tests. Things look cleaner than I thought they might. Due to a lack of test boxes, we'll have to do the Serial Command stuff on Friday.
    • 8/1: Only Mike and Bob showed up for testing today, but -- as luck would have it -- we discovered a procedural glitch in the way the 'scope had been set up yesterday. The net effect was that the sampling bandwidth was too slow and filtering out all the transition glitches. When we went for full bandwidth, all the dirt showed up.
      Not surprisingly, since DAG did not put the speed-ups in her circuitry, the DPA telemetry does not show leading edge ringing.
    • 8/2: Mike put a tap onto the Serial Cmd Clock in the 6" jumper we have between the Flight Cable and the DPA ('cause the Flight Cable can't be plugged into the DPA). The leading edge ripple is still there, although looks best with an ungrounded 'scope(!!).
    • 8/5: TRW tried an experiment with just a 300 ohm resistor on the bulkhead connector; with the 'scope speed slow enough, there was no problem (they evidently didn't look at faster time scales!).
    • 8/12: TRW showed up, but Gulton did not, for the grand serial digital test. MSFC in the person of Mike Goode got lost. The glitch is there, even when rerunning the "300 ohms at the bare bulkhead connector" test from last week! Also reviewed the HRC data wherein they reported "no problem". The "problem" with the data was obvious; they had their scope bandwidth limited to 20MHz (and Scott and Barry from TRW agree with our interpretation of the HRC data).
    • 9/15: ACIS, HRC, TRW, and BASD worked up a plausible schedule for running Science/Software/Hardware tests during the TB/TV testing. BASD continues to be reluctant to do full functional testing of the SIM, and BASD tiger teams are working the motor-chatter problem. Good news/bad news: he engineering hardware chatters just like the flight hardware.
    • 9/18-22: With the usual confusion and hurry-up-and-wait, ACIS got mechanically integrated -- except the radiator and shades, which we are keeping clean in a box.
    • 9/26: HRC had problems running, though they got some testing done with the SIM components. Then ACIS had general noise (traced to a cabling problem with the CTUE). With that fixed, found that the High Speed Tap was non-functional. Looks like a problem with cabling from the instrument to the buffer box.
      No crosstalk noise was noted, but three "liens" on the system exist: the HST is functioning, the PSMC lid temperature is not being reported properly in the data stream, the hardware serial commands aren't being accepted by the CTUE.
    • 10/21: CCD dark current data taken at -20, -40, -60, -90C as instrument did its cooldown. The DA took 24hr to reach -60 (!) and the FP bottomed out at -110C.
    • 10/22: The data displays have been updated with certain ISIM temperatures; after some fudging, the values agree with the Ball data logging programs to within a degree C.
      Since HRC needed Format 0 for 6 hours during initial HV turnon, we started in 500bps mode and will get an extra 6 hours in phase 3a.
    • 10/24: Cold, for the ISIM, is really cold. The therms are considering if it be wise to even go into the survival mode.
    • 10/26: Doing cold survival and running ahead of schedule. With the therms acting like engineers (and not representatives of their various organizations) things progress apace.
    • 10/27: It took 10 hours to get SIM mechanism up to cold operating from cold survival, then 4 hours to do a SIM functional.
    • 10/29: Visual inspection of the SIM table showed much mechanical grief in the moving parts. Disassembly is required.
    • 1/10: Unspecified difficulties in running SIM functional tests extended the planned 2hr test to 16hr. Then when the chamber shrouds were purged the entire chamber vented up to 0.2 torr before control was regained. A good thing our Detector Assembly was sealed.

      Test Notes from Science

      • Musings from Tuesday, 10/22
      • Musings from Wednesday, 10/23
      • Musings from Thursday, 10/24

        Anomalies, Major and Minor:

        • Given a loose connector in the diode-combiner box which sits between the DPA and the LRCTU, we managed to burn garbage into both BEPs -- with flakey serial data connections, much is possible. Fixed that problem and burned Flight 1.5 just fine.
        • The serial RCTU clocks (for some reason, possibly just test setup, the data bits aren't as bad) have ringing on their leading edges -- and some lesser amounts on their trailing edges; magnitude of the ring is typically just under a volt. This is bad.
        • The RCTU puts out race glitches on the data lines, up to a couple of volts. This isn't nice, but the data is stable at the sampling times.
        • The Focal Plane only got to -113 unregulated in the hot operating condition and -120 in the cold, 2CCD mode. This isn't cold enough by at least 10C. No substantive changes occured during retest in TV.

          Phone Numbers

          BASD GSE area number is: 303-939-6589
          BASD 2nd floor office: 303-939-5114
          BASD 2nd floor FAX: 303-939-5640
          Brutus ACIS location: 303-939-5431
          Brutus HRC location: 303-939-6814
          Brutus CTUE location: 303-939-6812

          MIT Boulder Field Center: 1853 26th Street, #309, Boulder, CO 80302; 303-541-9145

          Name Function Work Home Notes
          Demitrios Athens Software GSE 258-5285 617-566-7668 Ski Bum: 12/24->1/4
          Vacation: 1/8->1/14
          Mark Bautz Proj. Science 253-0023 617-863-2024
          Ed Boughan System Eng. 253-6146 508-470-1043 pager: 800-439-9827
          Mike Doucette Electrical GSE 253-3571 617-438-7690
          Peter Ford Software 253-7277 617-661-0868
          Bob Goeke Project Eng. 253-1910 617-729-8484 pager: 800-453-3658
          Gordon Gong Analog 253-8135 617-969-3768
          Dorothy Gordon Digital 253-0797 617-232-7327
          Fred Kasparian Mechanical 253-1065 617-327-6330
          Brian Klatt Quality 253-7555 508-785-1511
          Bill Mayer Project Mgr. 253-7552 508-692-3318
          Jim O'Connor Manufacturing 253-6139 617-335-2915
          Ellen Sen Thermal 253-8342 617-665-3568
          Matt Smith Contamination 252-1736 508-435-9970
          Mark Bautz Project Science 253-0023 617-566-6320
          Beverly LaMarr Science Coordinator 258-8153 617-776-0621 Lab: 3-7466
          Vacation: 12/24->12/28
          Bart Graham MSFC Rep. 303-939-5105

          Jim Esaker BASD Liason 303-939-5655
          pager: 303-826-3935
          Dale McConnell
          303-939-4211
          pager: 303-890-2811


          Phase I T/V test record is still available.

          Phase II T/V test record is still available.

          Page maintainer: Bob Goeke.

          There have been 5 accesses to this page since 9/14/2022.